# Adapting Vision Foundation Models for Robust Cloud Segmentation in Remote Sensing Images

### v1

@@ IV-C7 Overall Components in the Cloud-Adapter
In the baseline model, without any components from the spatial perception module or adapting module, the features from the VFM backbone were directly passed to the head for segmentation prediction. This configuration achieved a mIoU of 73.10%. When the stem module was added, the mIoU increased slightly to 73.12%, with a negligible parameter increase ($3\times 10^{-4}$). Including the stem module helps in early spatial feature projection and extraction, providing the model with a rudimentary spatial awareness information.
@@
Finally, with the inclusion of the adapting module, the full model achieved the highest mIoU of 74.18% and a total of 1.82M parameters. The adapting module enables fine-tuned integration of the spatial and backbone features, significantly enhancing the model’s performance despite increased parameters. This demonstrates the crucial role of the adapting module in optimizing the feature fusion process. In conclusion, the ablation study confirms that each component contributes incrementally to the model’s performance. The concatenation of spatial and backbone features, followed by successive refinement and adaptation through the added modules, substantially improved the segmentation performance in special domains.
@@ IV-D1 Binary-Class Cloud Segmentation
For binary-class cloud segmentation datasets, Cloud-Adapter consistently demonstrates superior performance across all evaluation metrics (aAcc, mIoU, mAcc, and mDice). On the HRC_WHU dataset (Table VII), Cloud-Adapter achieves the highest scores with an aAcc of 94.50%, mIoU of 89.05%, mAcc of 93.74%, and mDice of 94.19%, significantly outperforming the second-best method, HRCloudNet, which only achieves a mIoU of 83.44%. On the GF12MS_WHU_GF1 dataset (Table VIII), Cloud-Adapter attains an aAcc of 98.92%, mIoU of 92.55%, mAcc of 95.99%, and mDice of 96.02%, surpassing KappaMask, which achieved a slightly lower mIoU of 92.42%. On the GF12MS_WHU_GF2 dataset (Table IX), Cloud-Adapter maintains its leading performance with an aAcc of 94.04%, mIoU of 83.02%, mAcc of 87.54%, and mDice of 90.40%, showing its robustness in handling challenging satellite data.
@@
Fig. 6 compares the visualized segmentation results for three binary-class datasets. Cloud-Adapter stands out for its ability to accurately capture cloud boundaries and reduce over-segmentation issues, which are common in other methods like DBNet and HRCloudNet. The results clearly illustrate its superiority in segmenting clear sky and cloud regions, even under complex conditions.
@@ IV-D2 Multi-Class Cloud Segmentation
Cloud-Adapter (Ours) 88.99 55.04 85.31 65.64 73.75 94.02 74.93 91.48 77.41 84.46 94.17 71.00 92.07 79.26 84.13 89.99
@@
Cloud-Adapter achieves state-of-the-art performance on multi-class cloud segmentation datasets, which include four categories: Clear Sky (CRS), Thin Clouds (TNC), Thick Clouds (TKC), and Cloud Shadows (CDS). On the CloudSEN12_High_L1C dataset (Table X), Cloud-Adapter achieves a mIoU of 74.18%, with IoUs of 89.19% (CRS), 85.46% (TKC), 56.15% (TNC), and 65.93% (CDS). It outperforms methods like UNetMobv2 and HRCloudNet, which show lower scores, particularly in challenging classes such as TNC and CDS. On the CloudSEN12_High_L2A dataset (Table XI), Cloud-Adapter achieves a mIoU of 73.75%, with IoUs of 88.99% (CRS), 85.31% (TKC), 55.04% (TNC), and 65.64% (CDS). Its mDice score of 84.13% further highlights its balanced performance across all categories. On the L8_Biome dataset (Tables XII and XIII), Cloud-Adapter attains the highest mIoU (57.53%), mAcc (69.34%), mDice (70.46%), and aAcc (81.79%). It excels in challenging categories like Snow/Ice (32.01%) and Shrubland (51.18%), consistently outperforming competitors such as DBNet and UNetMobv2 across diverse land scenes.
@@
Fig. 7 and 8 present Cloud-Adapter’s segmentation results on multi-class datasets (CloudSEN12_High_L1C and CloudSEN12_High_L2A), accurately distinguishing thick clouds, thin clouds, and cloud shadows in complex conditions. The figures also demonstrate its performance across varying cloud cover levels: “Low” (less than 35%), “Medium” (35–65%), and “High” (over 65%). These results highlight Cloud-Adapter’s robustness and adaptability to diverse atmospheric scenarios. Fig. 9 highlights Cloud-Adapter’s segmentation performance on the L8_Biome dataset, demonstrating superior generalization across diverse biomes, such as barren, forest, and snow/ice regions. These results further emphasize the model’s capability to adapt to varying environmental conditions while maintaining high accuracy and precision, solidifying its effectiveness for complex cloud segmentation tasks in remote sensing applications.
In conclusion, across both binary-class and multi-class datasets, Cloud-Adapter consistently outperforms state-of-the-art methods. Its high aAcc, mIoU, mAcc, and mDice scores demonstrate its capability to effectively handle diverse cloud segmentation tasks. The visual comparisons further validate its robustness and adaptability, making it a reliable algorithm for cloud detection in remote sensing applications.
@@ V Conclusion
Limitations. The performance of Cloud-Adapter depends heavily on the VFM, and the large parameters of VFMs make it difficult to deploy on edge devices with limited resources.

### v2

@@ IV-C7 Overall Components in the Cloud-Adapter
In the baseline model, without any components from the spatial perception module or adapting module, the features from the VFM backbone were directly passed to the head for segmentation prediction. This configuration achieved a mIoU of 73.10%. When the stem module was added, the mIoU increased slightly to 73.12%, with a negligible parameter increase ($3\times 10^{-4}$). Including the stem module helps in early spatial feature projection and extraction, providing the model with rudimentary spatial awareness information.
@@
Finally, with the inclusion of the adapting module, the full model achieved the highest mIoU of 74.18% and a total of 1.82M parameters. The adapting module enables fine-tuned integration of the spatial and backbone features, significantly enhancing the model’s performance despite increased parameters. This demonstrates the crucial role of the adapting module in optimizing the feature fusion process. In conclusion, the ablation study confirms that each component contributes incrementally to the model’s performance. The concatenation of spatial and backbone features and successive refinement and adaptation through the added modules substantially improved the segmentation performance in particular domains.
@@ IV-D1 Binary-Class Cloud Segmentation
Cloud-Adapter consistently demonstrates superior performance across all evaluation metrics (aAcc, mIoU, mAcc, and mDice) for binary-class cloud segmentation datasets. On the HRC_WHU dataset (Table VII), Cloud-Adapter achieves the highest scores with an aAcc of 94.50%, mIoU of 89.05%, mAcc of 93.74%, and mDice of 94.19%, significantly outperforming the second-best method, HRCloudNet, which only achieves a mIoU of 83.44%. On the GF12MS_WHU_GF1 dataset (Table VIII), Cloud-Adapter attains an aAcc of 98.92%, mIoU of 92.55%, mAcc of 95.99%, and mDice of 96.02%, surpassing KappaMask, which achieved a slightly lower mIoU of 92.42%. On the GF12MS_WHU_GF2 dataset (Table IX), Cloud-Adapter maintains its leading performance with an aAcc of 94.04%, mIoU of 83.02%, mAcc of 87.54%, and mDice of 90.40%, showing its robustness in handling challenging satellite data.
@@
Fig. 6 compares the visualized segmentation results for three binary-class datasets. Cloud-Adapter stands out for its ability to accurately capture cloud boundaries and reduce over-segmentation issues, which is common in other methods like DBNet and HRCloudNet. The results illustrate its superiority in segmenting clear sky and cloud regions, even under complex conditions.
@@ IV-D2 Multi-Class Cloud Segmentation
Cloud-Adapter (Ours) 89.04 54.97 84.91 64.60 73.38 94.65 73.64 90.61 76.81 83.93 94.20 70.95 91.84 78.49 83.87 89.90
@@
Cloud-Adapter achieves state-of-the-art performance on multi-class cloud segmentation datasets, which include four categories: Clear Sky (CRS), Thin Clouds (TNC), Thick Clouds (TKC), and Cloud Shadows (CDS). On the CloudSEN12_High_L1C dataset (Table X), Cloud-Adapter achieves a mIoU of 74.18%, with IoUs of 89.19% (CRS), 85.46% (TKC), 56.15% (TNC), and 65.93% (CDS). It outperforms methods like UNetMobv2 and HRCloudNet, which show lower scores, particularly in challenging classes such as TNC and CDS. On the CloudSEN12_High_L2A dataset (Table XI), Cloud-Adapter achieves a mIoU of 73.38%, with IoUs of 89.04% (CRS), 84.91% (TKC), 54.97% (TNC), and 64.60% (CDS). Its mDice score of 83.87% further highlights its balanced performance across all categories. On the L8_Biome dataset (Tables XII and XIII), Cloud-Adapter attains the highest mIoU (57.53%), mAcc (69.34%), mDice (70.46%), and aAcc (81.79%). It excels in challenging categories like Snow/Ice (32.01%) and Shrubland (51.18%), consistently outperforming competitors such as DBNet and UNetMobv2 across diverse land scenes.
@@
Fig. 7 and 8 present Cloud-Adapter’s segmentation results on multi-class datasets (CloudSEN12_High_L1C and CloudSEN12_High_L2A), accurately distinguishing thick clouds, thin clouds, and cloud shadows in complex conditions. The figures also demonstrate its performance across varying cloud cover levels: “Low” (less than 35%), “Medium” (35–65%), and “High” (over 65%). These results highlight Cloud-Adapter’s robustness and adaptability to diverse atmospheric scenarios. Fig. 9 highlights Cloud-Adapter’s segmentation performance on the L8_Biome dataset, demonstrating superior generalization across diverse biomes, such as barren, forest, and snow/ice regions. These results further emphasize the model’s capability to adapt to varying environmental conditions while maintaining high accuracy and precision and solidify its effectiveness for complex cloud segmentation tasks in remote sensing applications.
In conclusion, Cloud-Adapter consistently outperforms state-of-the-art methods across binary and multi-class datasets, demonstrating its ability to effectively handle diverse cloud segmentation tasks. The visual comparisons further validate its robustness and adaptability, making it a reliable algorithm.
@@ V Conclusion
Limitations. The performance of Cloud-Adapter depends heavily on the VFM, and the large parameters of VFMs make it difficult to deploy on edge devices with limited resources. Additionally, in the complex dataset (L8_Biome), which encompasses diverse land cover scenarios, our method achieved a leading mIoU score of 57.53%. However, a substantial gap remains when compared to manually annotated masks. This highlights the potential for improving algorithm efficiency and developing new techniques to enhance segmentation accuracy.